# Automated Multi-Task Learning for Joint Disease Prediction on Electronic Health Records

### v1

@@ 4.1.3. Evaluation Metric
Table 2. Hyperparameter setting.
@@
Figure 2. Histogram of task gains for AutoDP in terms of Averaged Precision.
4.1.4. Implementation
We implement the framework using the PyTorch framework and run it on an NVIDIA A100 GPU. Given the dataset we have, we first train a vanilla LSTM for every task independently, and report the backbone performance in Table 1, which can be further used to compute multi-task gains. For, the proposed method, we run three settings of experiments: Task @ 5, Task @ 10 and Task @ 25, which refers to using the first 5 tasks, 10 tasks and 25 tasks respectively. For different settings, we use specific hyperparameters as shown in Table 2. Besides that, we define the candidate operation set $\mathcal{O}$ as {Identity, Zero, FFN, RNN , Attention}, which includes widely used operations for processing EHR time series. Among them, Identity means maintaining the output identical to the input. Zero means setting all the values of the input feature to 0. Attention and FFN represents one self-attention layer and one feed-forward layer respectively, which are the same as in Transformer (Vaswani et al., 2017). RNN is one recurrent layer, and we adopt LSTM (Hochreiter and Schmidhuber, 1997) in our framework. For all the MTL procedures and baseline training, we apply the batch size of 64 and learning rate of $43e-$ 4. For training the surrogate model, we use the batch size of 5 and learning rate of $55e-$ 5. During searching, we compute all multi-task gains on the validation set for guiding the surrogate model training. After we obtain the optimal configuration, we train the searched models and report their multi-task gains on the test set.
Table 3. Performance comparison in terms of averaged per-task gain over single task backbone (All results are in the form of percentage values %).
Task Groups Tasks @ 5 Tasks @ 10 Tasks @ 25
Metric ROC AVP ROC AVP ROC AVP
LSTM +0.09 +0.18 +1.06 +3.22 +1.83 +7.46
Transformer +0.97 +4.82 +1.41 +4.14 +1.75 +7.45
Retain +0.46 +1.80 +0.66 +0.75 +1.41 +5.88
Adacare +1.03 +5.21 +1.32 +4.05 +1.68 +6.94
DARTS +1.28 +5.01 +2.01 +6.87 +1.87 +7.71
MTG+LSTM +0.51 +2.10 +0.65 1.87 +1.74 +7.40
MTG+Transformer +0.91 +3.64 +1.20 +3.95 +1.79 +9.15
MTG+Retain +0.55 +3.11 +1.51 +5.20 +1.54 +8.87
MTG+Adacare +1.25 +5.78 +1.44 +4.63 +1.75 +7.84
MTG+DARTS +1.47 +6.41 +2.02 +6.65 +2.41 +11.76
AutoDP +1.49 +7.12 +2.08 +7.53 +2.68 +12.70
4.2. Performance Evaluation
We compare our method with various baselines and show the results in Table 3. Our method can outperform the baselines at all three settings in terms of averaged per-task gain for ROC and AVP, which is a significant improvement over existing MTL frameworks for EHR data.
First, without considering task grouping, we train one shared model to predict for all tasks in three settings and compute the multi-task gain for them. For the hand-crafted encoders, they can only bring minimal average performance gain over single-task backbones, especially for Task @ 5. Note that even if the average per-task gain is positive, there is a great chance for some tasks to have negative gains. Also, we apply DARTS to the architecture search space we defined $\mathcal{A}$ for each setting and use the searched encoder to predict for all tasks. We can observe that the searched architectures generally perform better than hand-crafted models, but might get worse results in some rare cases. Without the task grouping, it is difficult to achieve higher performance gains even for searched architectures. This means naively putting all tasks together to train will not provide much improvement over single task training.
Moreover, we combine all the encoders with the multi-task grouping baseline to achieve better results. Specifically, we first run the MTG method to perform task grouping under the same budgets with our method as defined in Table 2. For each task group, we employ all four EHR encoders and the DARTS search algorithm. The results show that the sequential optimization over task grouping and architecture search (MTG+DARTS) performs better than other hand-crafted encoders. Compared to the best baseline, our method is able to achieve higher gains at all three setting. We attribute this to the superiority of joint optimization over task grouping and architecture search, which can potentially find global optimal configurations for both aspects.
@@ 4.2. Performance Evaluation
4.3. Hyperparameter Analysis
Here, we analyze the effect of two vital hyperparameters of our method: $K_{1}$ and $B$. We try out different values of them under the setting of Task @ 25 and report the corresponding averaged per-task gain of AVP in Figure 3.
First, $K_{1}$ determines the number of training samples collected during searching. Given that each sample involves an MTL procedure, it constitutes a major portion of the computation cost. Therefore, our goal is to find an optimal value for $K_{1}$, striking a balance between cost-effectiveness and achieving commendable performance. Examining the left figure, we observe a direct correlation between the increase of $K_{1}$ and a corresponding improvement in performance. This relationship is obvious, as more rounds lead to more training data for the surrogate model. Furthermore, we notice that the performance change become less pronounced after $K_{1}$ reaches 25. This suggests that the surrogate model effectively learns the distribution of the search space after consuming $25\times 25$ training samples during active selection (25 samples per round). As a result, we can empirically decide to halt the iteration at this point with little impact on performance.
Second, $B$ determines the number of task groups for the final configuration, which indicates the number of MTL models needed for achieving the expected performance gain after searching. We also observe similar phenomenon that the performance becomes stable after $B$ reaches certain amount. We could also choose the optimal value for $B$ accordingly.
Figure 3. Analysis for the number of active sampling rounds $K_{1}$ and the budget of task groups $B$ under the setting of Task @ 25.
Figure 4. Comparing the proposed active sampling strategy with random sampling.
4.4. Random vs Active Sampling
We further analyze the effect of the proposed active sampling strategy. In order to demonstrate its effectiveness, we conduct a comparative analysis with the random sampling method. In the case of the random method, an equivalent number of training samples is randomly selected from the search space to train the surrogate model. Then, the trained surrogate model is employed for deriving the final configuration in same way as introduced in Section 3.5. Finally, we compare the averaged per-task gain of AVP achieved by each approach, which are shown in Figure 4.
Under all three settings, the proposed active sampling strategy is able to achieve significant improvement over random baseline, ranging from $5.48\%$ to $12.39\%$. It is noteworthy that active sampling tends to have larger improvement when the number of tasks grows. We attribute this to the expansion of the search space. By progressively selecting samples capable of yielding higher performance gains during searching, the active sampling method has superior sample efficiency than the random method. Therefore, in the case of Task @ 25, where the fraction of samples with ground-truth evaluations is exceptionally small (approximately $2e-5\%$), the active sampling method can achieve larger improvement over random method.
Figure 5. Illustration of the searched configuration under the setting of Task @ 10.
4.5. Visualization of the Searched Configurations
@@
Here, we show one example of the final configuration for setting Task @ 10 in Figure 5. The proposed AutoDP identifies 5 different task groups and also searches for the corresponding architectures. We can observe that some of the tasks tend to be trained independently, while others are grouped together for joint training. This supports our claim that fine-grained task grouping is necessary to bring the optimal performance gain. Also, the optimal architecture is also different for each task group, which further justifies the necessity of joint optimization over task grouping and architecture search.
5. Conclusions
In this paper, we propose AutoDP, an automated multi-task learning framework for joint disease prediction on EHR data. Compared to existing work, our method largely improves the design of task grouping and model architectures by reducing human interventions. Experimental results on real-world EHR data demonstrate that the proposed framework not only outperforms existing state-of-the-art methods, but also maintains a feasible search cost. In future work, we will try to enable the searching for more aspects of MTL, such as hyperparameters, optimization methods, etc.

### v2

@@ Appendix A Implementation
Table 4: Hyperparameter setting.
@@
Appendix B Individual task gains
@@
Figure 3: Histogram of task gains for AutoDP in terms of Averaged Precision.
Appendix C Greedy search
Here, we show the detailed procedure of the greedy search algorithm for deriving the final MTL configuration in Algorithm 2. Given the budget $B$, we aim at searching for $B$ samples from the search space $\mathcal{C}\times\mathcal{A}$ such that the overall gain $G$ is maximized. Specifically, we first randomly initialize the population $\mathcal{P}$ that contains $B$ pairs of task combinations and architectures. Then, at every iteration, we randomly mutate one pair $(C,A)$ from the population and see whether the overall multi-task gain will increase. If so, we update $\mathcal{P}$ accordingly. After $K_{2}$ iterations, we can obtain a near-optimal solution. In practice, we also apply multiple initial populations to avoid getting stuck on local optima. Although we only get an approximate solution, our method can already achieve significant improvements over baselines as shown in Section 4.2.
1
Input: Trained surrogate model $F(\cdot)$, $B$, $K_{2}$.
Output: Searched population $\mathcal{P}$.
3 for $v=1,2,\cdots,K_{2}$ do
4       Randomly select one pair $(C,A)$ from $\mathcal{P}$;
6       Estimate $\mathcal{P}^{\prime}$ and $\mathcal{P}$ using $F(\cdot)$;
8
9 end for
10
Algorithm 2 Greedy Search
Appendix D Disease Based Grouping
To show the effectiveness of automated task grouping, we conduct experiments using a predefined task grouping based on disease categories. We asked GPT-4 [34] to classify the 25 prediction tasks into different groups based on their medical meaning. The result is shown in Table 5. Using this grouping, we further apply DARTS to each group and report the multi-task gains as shown in Table 2. Compared to AutoDP, there is a notable performance drop for the disease based grouping. This means human intuition dose not provide the optimal task grouping, which underscores the necessity of employing search algorithm to automatically discover better task grouping for MTL.
Table 5: Disease Based Grouping.
Groups Diseases
Cardiovascular Diseases
Acute cerebrovascular disease
Acute myocardial infarction
Cardiac dysrhythmias
Congestive heart failure; nonhypertensive
Coronary atherosclerosis and other heart disease
Essential hypertension
Respiratory Diseases
Chronic obstructive pulmonary disease and bronchiectasis
Other lower respiratory disease
Other upper respiratory disease
Pleurisy; pneumothorax; pulmonary collapse
Pneumonia (except that caused by tuberculosis or sexually transmitted disease)
Respiratory failure; insufficiency; arrest (adult)
Kidney Diseases
Acute and unspecified renal failure
Chronic kidney disease
Metabolic Diseases
Diabetes mellitus with complications
Diabetes mellitus without complication
Disorders of lipid metabolism
Fluid and electrolyte disorders
Gastrointestinal Diseases
Gastrointestinal hemorrhage
Infections
Septicemia (except in labor)
Surgical/Medical Complications
Complications of surgical procedures or medical care
Neurological/Cardiac Conditions
Conduction disorders
Shock
Liver Diseases
Other liver diseases
Appendix E Visualization of the Searched Configurations
@@
Here, we show two example of the final configuration for setting Task @ 10 in Figure 4 and for Task @ 25 in Figure 5. The proposed AutoDP identifies 5 and 10 different task groups respectively and also searches for the corresponding architectures. We can observe that some of the tasks tend to be trained independently, while others are grouped together for joint training. This supports our claim that fine-grained task grouping is necessary to bring the optimal performance gain. Also, the optimal architecture is also different for each task group, which further justifies the necessity of joint optimization over task grouping and architecture search.
Figure 4: Illustration of the searched configuration under the setting of Task @ 10.
Figure 5: Illustration of the searched configuration under the setting of Task @ 25.
Appendix F Limitations and future work
There are some valuable future directions based on the current version of AutoDP.
First, from the application perspective, if we aim at deploying AutoDP to real-world healthcare systems, it would be advantageous to apply it to more complex problem settings. For example, the incorporation of diverse clinical data sources beyond EHR such as claims, drugs, medical images and texts will significantly enhance the practical utility of AutoDP. Additionally, considering the dynamic nature of healthcare environments with continuously updated input data and evolving tasks, adapting the surrogate model to accommodate new data and tasks would be imperative.
Second, addressing privacy concerns within healthcare systems is a promising direction. Therefore, extending AutoDP with data processing pipelines for automatic feature engineering could offer enhanced privacy safeguards and further improve its applicability in sensitive healthcare contexts.